# Language Modeling for Fast Inference

### v1

@@ 3.4 Ablation on components of the Block Transformer
3(f) shows the training loss curve for three token decoder strategies. Using a cross-attention module with key and value sequences equal to the block length considerably diminishes performance. In contrast, forwarding context embeddings through self-attention operations enhances performance, with prefix decoding surpassing other methods. Furthermore, extending the prefix beyond four tokens markedly improves perplexity, effectively broadening the computation width of token decoder. Since longer prefixes add minimal inference overhead, we select a prefix length of two by balancing performance with FLOPs. This approach offers new insights into global-to-local modeling, diverging from previous studies [74] which overlook the potential of local computational capacity in the token decoder. Detailed results across various model sizes are summarized in Appendix M.
@@ 3.5 Analysis on global-to-local language modeling
Since the token decoder depends solely on the context embedding, there could be a concern about whether the Block Transformer fully utilize context information. To address this, we evaluate the loss of token positions within a 2K context window using the test set of PG19 dataset [52]. 4(b) indicates that later tokens are consistently predicted with higher likelihood, suggesting that our architecture, which distinguishes between block-level and token-level decoders, effectively leverages at least 2K tokens of context.
@@ 3.7 Uptraining from vanilla transformers
Unlike previous studies [74], our subword-level global-to-local architecture can leverage the initialization from a pretrained vanilla transformer. This enables efficient training, requiring only a small number of data. As shown in 5(a), this uptraining strategy can lead to near-full performance recovery with just 10% of the original training steps, outperforming random initialization strategy. Consistent with previous studies [2], investigating deliberate weight initialization techniques can further enhance the performance convergence. We summarize details in Appendix N.
4 Discussion
@@ 4.1 Comparison to related works
The MEGABYTE model [74] adopts a global-to-local structure but focuses on efficient pretraining over inference. Thus, within the training FLOPs budget, they argue for a larger block decoder based on a 6:1 ratio deemed optimal. As shown in 5(b), we reimplement the token-level MEGABYTE models, and they also achieve significantly higher throughput compared to vanilla models through global-to-local modeling. Nevertheless, consistent with our insights in subsection 3.3, our models with enhanced local computational capacity demonstrate a significant throughput increase of over 1.5 times on top of MEGABYTE. See Appendix O for more details.
@@
Global-to-local modeling can be viewed through the lens of KV cache compression, where past sequences are entirely pruned in the upper layers. Recent studies have introduced algorithms that preserve only meaningful tokens, determined by accumulated attention scores [67, 77], with observing that most attention tends to sink into the first token [72, 28]. In 5(c), our models exhibit a similar pattern. This observation suggests that performance could be enhanced by leveraging not just the current context embedding but also by incorporating global embeddings or context embeddings from the previous window. See Appendix P for more details.
@@ 4.2 Contextual information encapsulated in context block embedding
Since the input tokens and context embeddings share the same latent space in the token decoder, we analyze the nearest tokens to these block embeddings. Interestingly, Table 5 in Appendix Q reveals that context embeddings compress global context rather than outlining the next block. The second prefix often contains information about the last token of current block to aid predicting the first token of the next block. Meanwhile, the first prefix typically matches non-intuitive or the EOS token, suggesting that they carry more general information. In light of this, the block decoder effectively compresses past global contexts, which the token decoder leverages for its local language modeling.
@@ 5 Conclusion
We would like to thank Honglak Lee for his critical feedback on the outline of the paper. We also extend our gratitude to Yujin Kim for extensive discussions on efficient inference and related work. Additionally, we appreciate the continuous feedback from Kyungmin Lee, Junwon Hwang, Park Sangha, and Hyojin Jeon, throughout the development of our work.
Appendix A Limitation
@@
Appendix B Future works
@@ Appendix C Broader impact
Appendix D Related work
D.1 Global-to-local modeling
While prior research has explored global-to-local modeling in various modalities, it has not been utilized in autoregressive language models to optimize inference efficiency. Local modules have been incorporated in vision transformers to enhance the encoding of fine-grained local features with minimal computational overhead [31]. A similar approach has been used in BERT-based encoder LMs to effectively classify long documents by first encoding each fixed-size segments independently [49, 19, 43]. This is similar to how we put emphasis on the local token decoder, but we apply locality to decoder LMs at the upper layers rather than encoder LMs at lower layers.
Byte-level hierarchical models
Several works on byte-level modeling apply a similar architecture to our Block Transformer [74, 44]. However, while we attempt to mitigate the bottlenecks of global attention by isolating the task of language modeling into global and local components, prior work mainly utilize the hierarchical structure to mitigate the long contexts lengths of byte-level data from various modalities in the absence of tokenization. In contrast to the central role of the local module in our Block Transformer, prior work consider the role of the local model as ‘mapping a hidden state to a distribution over possible patches’, and suggest that ‘much smaller model can be used for inter-patch modelling’ [74] and may ‘cease to contribute to overall performance’ [44]. Similarly, while Yu et al. [74] finds that it is optimal to assign more parameters to the global module under same training-time constraint, we find that a more balanced allocation, e.g., 1:1 for $L_{B}=4$, is optimal under fixed parameter constraints, and that even larger token decoders are beneficial for inference throughput, further highlighting the benefits of the local module. We believe these differences in interpretation and findings stem from a difference in the granularity of input units, i.e., bytes vs subwords, as well as the cost metrics considered, i.e., training-time vs inference cost.
D.2 KV cache compression
@@
D.3 Architectures for optimizing KV cache
@@ Appendix H Random length padding during pre-training
To apply inference on prompts whose lengths are not multiples of $L_{B}$, we need to add padding tokens to the prompt to fill the input blocks. Unlike padding tokens in vanilla transformers, these padding tokens are actually considered in the computation of the input block embedding, due to the fixed-size nature of our embedding methods, except for the CLS token variant. Therefore, we add random padding tokens with uniform length between 0 and $L_{B}-1$ at the beginning of each document when applying input packing during pre-training. We also pad the unfilled tokens in the last block of each document, to prevent multiple documents being included in a single block. Note that this was applied after our main experiments, thus were not applied to our largest models in Table 2. We posit that this has adversely affected some downstream task performance evaluations.
@@
Appendix I Pareto frontiers at variable batch sizes and context lengths
@@
Appendix J Position-wise loss by parameter allocation ratio
@@
Appendix K Loss trend by allocation ratio and block length
@@
Appendix L Pareto frontier of throughput by allocation ratio and block length
@@
Appendix M Ablation studies on components of Block Transformer
M.1 Embedder design
@@
M.2 Token decoder design
@@
Appendix N Uptraining strategy for training efficiency
@@
Ainslie et al. [2] have demonstrated the significance of weight initialization for effectively uptraining models. Our extensive ablation studies reveal the optimal strategies for the Block Tramsformers: (1) Dividing a vanilla transformer layer in half and assigning each half to the block and token decoders, respectively, outperforms assigning the same weights of selected layers to both. (2) Initializing the input block embedding as the average of token embeddings within the block improves performance. (3) Initilizing token decoder prefixes by replicating the context embedding enhances convergence. As depicted in Figure 18, these initialization technqiues allow uptrained models to nearly match fully pretrained models. While larger models generally require longer uptraining, this approach still converges faster and recovers performance better than random intialization.
@@
Figure 18: Training loss curve of uptraining strategy for two different model sizes. Scratch denotes pretraining models from randomly initialized weights. The numbers in parentheses represents the number of training tokens.
Appendix O Performance comparison to MEGABYTE
@@
Appendix P Visualization of attention scores in Block Transformer
@@
Appendix Q Analysis on the context block embedding

### v2

@@ 3.4 Ablation on components of the Block Transformer
3(f) shows the training loss curve for three token decoder strategies. Using a cross-attention module with key and value sequences equal to the block length considerably diminishes performance. In contrast, forwarding context embeddings through self-attention operations enhances performance, with prefix decoding surpassing other methods. Furthermore, extending the prefix beyond four tokens markedly improves perplexity, effectively broadening the computation width of token decoder. Since longer prefixes add minimal inference overhead, we select a prefix length of two by balancing performance with FLOPs. This approach offers new insights into global-to-local modeling, diverging from previous studies [84] which overlook the potential of local computational capacity in the token decoder. Detailed results across various model sizes are summarized in Appendix N.
@@ 3.5 Analysis on global-to-local language modeling
Since the token decoder depends solely on the context embedding, there could be a concern about whether the Block Transformer fully utilize context information. To address this, we evaluate the loss of token positions within a 2K context window using the test set of PG19 dataset [62]. 4(b) indicates that later tokens are consistently predicted with higher likelihood, suggesting that our architecture, which distinguishes between block-level and token-level decoders, effectively leverages full context information. Furthermore, we observed the same trend with an 8K context length (in Figure 20), demonstrating our model’s ability to fully exploit at least 8K tokens of context. Our block language modeling further demonstrated its effective utilization of full context in the recent Needle-In-a-Haystack long-context benchmark [39] (refer to Appendix O for details.)
@@ 3.7 Uptraining from vanilla transformers
Unlike previous studies [84], our subword-level global-to-local architecture can leverage the initialization from a pretrained vanilla transformer. This enables efficient training, requiring only a small number of data. As shown in 5(a), this uptraining strategy can lead to near-full performance recovery with just 10% of the original training steps, outperforming random initialization strategy. Consistent with previous studies [2, 5], investigating deliberate weight initialization techniques can further enhance the performance convergence. We summarize details in Appendix P.
@@ 3.8 Comparison to related works
The MEGABYTE model [84] adopts a global-to-local structure but focuses on efficient pretraining over inference. Thus, within the training FLOPs budget, they argue for a larger block decoder based on a 6:1 ratio deemed optimal. As shown in 5(b), we reimplement the token-level MEGABYTE models, and they also achieve significantly higher throughput compared to vanilla models through global-to-local modeling. Nevertheless, consistent with our insights in subsection 3.3, our models with enhanced local computational capacity demonstrate a significant throughput increase of over 1.5 times on top of MEGABYTE. See Appendix Q for more details.
@@
Global-to-local modeling can be viewed through the lens of KV cache compression, where past sequences are entirely pruned (yet compressed into a context embedding) in the upper layers. Similar to techniques that preserve only meaningful tokens determined by accumulated attention scores [78, 87], this offers a promising way to improve decoding speed without compromising performance. Following the observations of [82, 32] that attention often concentrates on the first token (frequently semantically unimportant), 5(c) reveals a similar pattern in our block decoder. This suggests that augmenting the token decoder’s input with the initial “sink” block embedding, or a local window of embeddings, could yield substantial performance gains.
4 Discussion
@@ 4.1 Contextual information encapsulated in context block embedding
Since the input tokens and context embeddings share the same latent space in the token decoder, we analyze the nearest tokens to these block embeddings. Interestingly, Table 7 in Appendix S reveals that context embeddings compress global context rather than outlining the next block. The second prefix often contains information about the last token of current block to aid predicting the first token of the next block. Meanwhile, the first prefix typically matches non-intuitive or the EOS token, suggesting that they carry more general information. In light of this, the block decoder effectively compresses past global contexts, which the token decoder leverages for its local language modeling.
@@ 4.2 Techniques for further throughput improvement
5 Related work
The Block Transformer exemplifies an approach to coarse and local processing in autoregressive transformers. We compare our architecture with previous approaches for coarse and local processing, underscoring its unique effectiveness in alleviating bottlenecks in autoregressive inference.
Hierarchical transformers for coarse processing
Many previous works have explored hierarchical transformers to process long sequences efficiently. Early works use local encoders at early layers to obtain pooled representations of documents [59] or image patches [36]. Later works explore a downsample-then-upsample approach to process long sequences at coarser levels within the core of the model [22, 53]. This has been followed by numerous improvements including an autoregressive formulation [56], dynamic pooling lengths [57], and reduced decoding steps [28]. While this can reduce KV cache at middle layers, the upper and lower layers suffer from the same attention bottlenecks as vanilla transformer layers, where representations are at the finest level. In contrast, we identify these bottlenecks and mitigate them by applying local processing at the finer levels.
Local processing in modern language models
Early hierarchical transformers employ local computation at early layers [59, 36], but these are limited to encoder models. The most prominent adaptation of local processing in modern LMs is sliding window attention (SWA) [19, 8], adopted in GPT-3 [14] and Mistral [38]. While SWA reduces KV cache memory, window sizes are typically much longer than the block lengths used in Block Transformer; Mistral uses a window size of 2048, which requires 512 times more KV cache compared to our baseline token decoder. Previous adaptations of SWA typically incorporate global attention layers [14] or exploit stacked layers [38] to attend to older sequences. For example, with window size $W$, the current token attends to $W$ tokens of context in the final layer; these attend to up to $12W-$ 1 tokens in the previous layer; and so on. Due to this dependency, SWA does not benefit from the same prefill optimizations as the Block Transformer. Under our global-to-local approach, the token decoder restricts attention within block boundaries across all layers, enabling the prefill stage to be skipped for all preceding blocks.
Global-to-local hierarchical transformers
Several works on byte-level modeling employ a similar structure as our Block Transformer [84, 50]. However, while we aim to optimize autoregressive inference for subword-level LMs, prior work mainly utilize the hierarchical structure to mitigate the long context lengths of byte-level data in the absence of tokenization. Hence, in contrast to the central role of our local module (token decoder), prior work consider the role of their local module as ‘mapping a hidden state to a distribution over possible patches’, and suggest that ‘much smaller model can be used’ [84] and may ‘cease to contribute to overall performance’ [50]. Yu et al. [84] concludes that it is optimal to assign more parameters to the global module under fixed training-time constraint. In contrast, we find that a more balanced allocation, e.g., 1:1, performs better and faster under fixed parameter constraints, and that even larger token decoders can maximize inference throughput, showing that the local module can contribute to performance in an efficient manner. By recognizing the key bottlenecks in autoregressive inference, we believe our work presents a novel interpretation and uncovers previously unrecognized benefits of global-to-local hierarchical transformers.
@@ 6 Conclusion
We would like to thank Honglak Lee for his critical feedback on the outline of the paper. We extend our gratitude to Yujin Kim for extensive discussions on efficient inference and related work. Additionally, we appreciate the continuous feedback from Kyungmin Lee, Junwon Hwang, Sangha Park, and Hyojin Jeon, throughout the development of our work. Thanks to Joshua Ainslie for discussion during the early stages of development.
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.2019-0-00075, Artificial Intelligence Graduate School Program (KAIST), 10%) and the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. 2022-0-00871, Development of AI Autonomy and Knowledge Enhancement for AI Agent Collaboration, 90%).
Appendix A Limitations
@@
Appendix B Discussion and future works
@@ Appendix C Broader impact
Appendix D Extended related work
D.1 KV cache compression
@@
D.2 Architectural for optimizations of KV cache
@@ E.2 Inference-time advantages of block and token decoders
The following paragraphs provide a detailed presentation of the inference benefits associated with our proposed block and token decoder. For visual clarification, refer to Figure 6, and Table 2 offers a comparative analysis of actual computation speeds.
@@
Figure 6: Illustration of key advantages of Block Transformer over Vanilla Transformers. Each colored box represents a single input unit that is processed at each layer, and input tokens A to L are prompt tokens.(1) The local token decoder does not need to prefill the prompt, as it is not used in subsequent generation, whereas the upper vanilla layers require the KV values of all prompt tokens to generate token M and onwards. (2) The token decoder only needs to fetch KV cache from up to $L_{B}=4$ local tokens, while the upper vanilla layers needs to fetch from all previous tokens at each step, which can go up to the thousands or millions. (3) Since the block decoder operates at the block level, overall computation, memory overhead, and forward steps are reduced by $L_{B}=4$. KV cache IO is reduced quadratically. (4) Block transformer enables significantly higher batch size and thus overall higher compute utilization on identical hardware, as it only needs to preserve the KV cache of the blue and green parts in memory (green is minuscule for longer context lengths).
Table 2: Measurements on key advantages of Block Transformer during prefill and decode. Per-sample walltime of key operations at lower layers (block decoder) and upper layers (token decoder), for vanilla model with 300M non-embed params and a better performing Block Transformer with 1.2B non-embed params, using identical hardware (one H100 GPU). Refer to the caption of Figure 1 on the reason behind the significant walltime savings, despite using more parameters.
Vanilla T. Block T.
Prefill Decode Prefill Decode
Upper Layers Attention 19.94 56.21 N/A 2.41
(Token Decoder) FFN 0.86 1.42 N/A 0.55
Lower Layers Attention 19.94 56.21 1.96 5.00
(Block Decoder) FFN 0.86 1.42 0.68 0.09
(a) Prefill-heavy setting (2048/128)
Vanilla T. Block T.
Prefill Decode Prefill Decode
Upper Layers Attention 0.15 500.50 N/A 31.55
(Token Decoder) FFN 0.06 16.75 N/A 7.71
Lower Layers Attention 0.15 500.50 0.05 47.24
(Block Decoder) FFN 0.06 16.75 0.04 1.44
(b) Decode-heavy setting (128/2048)
@@ Appendix H Random length padding during pre-training
To apply inference on prompts whose lengths are not multiples of $L_{B}$, we need to add padding tokens to the prompt to fill the input blocks. Unlike padding tokens in vanilla transformers, these padding tokens are actually considered in the computation of the input block embedding, due to the fixed-size nature of our embedding methods, except for the CLS token variant. Therefore, we add random padding tokens with uniform length between 0 and $L_{B}-1$ at the beginning of each document when applying input packing during pre-training. We also pad the unfilled tokens in the last block of each document, to prevent multiple documents being included in a single block. Note that this was applied after our main experiments, thus were not applied to our largest models in Table 1. We posit that this has adversely affected some downstream task performance evaluations. Figure 7 presents a comprehensive overview of the results obtained with and without appending random padding during both training and inference stages.
@@
Appendix I Throughput Comparison with FlashDecoding
In modern LLM deployments, decoding speed enhancements through kernel fusion techniques like the FlashAttention algorithm [23] are generally employed. These mechanisms reduce the number of memory accesses, leading to faster decoding in LLMs. Consequently, the speed advantages of our Block Transformer, which minimizes KV cache size and memory accesses, could be a little diminished compared to a vanilla Transformer. To investigate this, we measured the maximum throughput with FlashDecoding applied, as illustrated in Figure 8. Interestingly, we observed an overall trend similar to that presented in Figure 2 in the main paper. Our model architecture still benefits significantly from FlashAttention for global attention within the block decoder, resulting in a considerable speed improvement of up to 31%.
Standard FlashDecoding
Models N-emb Prefill Decode Prefill Decode
Vanilla 5M 10.8 41.6 18.2 37.6
19M 6.9 19.1 10.2 20.2
85M 2.3 6.2 3.4 6.8
302M 0.8 2.1 1.3 2.4
Block 5M 272.3 809.5 396.2 623.4
19M 175.3 421.4 228.7 376.0
85M 59.0 134.7 76.3 127.0
302M 21.0 44.1 27.3 45.9
1.2B 12.4 25.7 13.5 24.4
(a) Prefill-heavy setting.
(b) Decode-heavy setting.
Figure 8: Pareto frontier of throughput to langauge modeling performance using FlashDecoding.
Appendix J Pareto frontiers at variable batch sizes and context lengths
@@ Appendix I Throughput Comparison with FlashDecoding
Appendix K Position-wise loss by parameter allocation ratio
@@
Appendix L Loss trend by allocation ratio and block length
@@
Appendix M Pareto frontier of throughput by allocation ratio and block length
@@
Appendix N Ablation studies on components of Block Transformer
N.1 Embedder design
@@
N.2 Token decoder design
@@
Appendix O Long-context modeling ability
8K context length on the PG19 dataset
To further support the effectiveness of our proposed block language modeling in capturing full context, we conducted experiments with an 8K context length (refer to Figure 20). However, due to limited computational resources, we pretrained only a 70M parameter vanilla model and a 170M parameter Block model. Following prior work [84, 82, 87] that used token position-wise perplexity on the PG19 dataset to demonstrate the utilization of global information in long contexts, we evaluated our Block Transformer in the same manner with 8K context length (refer to Figure 4(b) of the main paper for 2K context window). Even with a extended 8K context window, our models effectively utilized the full context, showing a decreasing loss trend as token position increased, similar to the vanilla model. Besides, consistent with Table 1 in the main paper, the 170M block model outperformed the 70M vanilla model in terms of perplexity.
Figure 20: Long-context modeling ability up to 8K tokens. Average loss at different token positions (512-token bins) within 8192-token snippets from the entire PG19 test set. Blue and green lines show Vanilla and Block Transformers, trained on 28.8B tokens with context length of 8192. Orange line shows a vanilla-loss-equivalent Block Transformer checkpoint, at 23.6B tokens. Both architectures achieve lower loss as context length increases, with the Block Transformer performing relatively stronger at early tokens. Block Transformer achieves lower loss at all positions up to 8K.
Performance on the Needle-In-a-Haystack task
To precisely evaluate long-context modeling ability of LLMs, recent benchmarks, such as Needle-In-a-Haystack (NIAH) [39], LongBench [7], and ZeroScrolls [69], are typically utilized. However, to the best of our knowledge, these benchmarks mostly evaluate instruction-tuned models, as opposed to our pretrained base models. Nevertheless, we slightly modified the tailored prompt with an instruction version to evaluate the performance. Following prior work [67], we construct the context by first sampling 2K-length snippets from concatenated essays written by Paul Graham as the “haystack”, and then inserting a “needle” containing key information in a random location. Following Reid et al. [67], we use this needle format: “The special magic {city} number is: {number}.” Here, {city} is a randomly chosen city name, and {number} is a random 7-digit number. We then append a prompt that queries to model to retrieve the 7-digit number. We consider two prompt formats:
  1. 1.
[…]
  2. 2.
[…]
    ”. Here, we used the exact same format as that in the needle to query the model.
We measured the accuracy by generating 20 new tokens, and considering a prediction correct if the generated text contains the 7-digit number. Tables 5 and 6 presents a comparison of accuracy between vanilla and Block Transformers across two different prompts. We found that Block Transformers perform equally or stronger than loss-equivalent vanilla models, consistently across (1) needle locations, (2) model scales and (3) prompt variants. These results confirm that the Block Transformer, like the vanilla models, can effectively retrieve global information contained within the 2K context length. With the Gemini prompt, we observed an accuracy trend that was very similar to the perplexity trend of the vanilla vs. block models. Near-perfect performance with the Verbatim prompt supports the long-sequence modeling capabilities of our models even when context information is squeeze into a single context embedding. We believe this parity between vanilla and Block Transformers on 2K context length will extend to 8K and beyond.
Table 5: Accuracy on Needle-In-a-Haystack task with Gemini prompt. Note that depth refers to the relative of the location of the needle within the haystack, in percentages.
Depth
Models N-Emb 0 10 20 30 40 50 60 70 80 90 100 Mean
Vanilla 19M    0.00%    0.00%    0.00%    0.00%    0.00%    0.00%    0.00%    0.00%    0.20%    0.80%    6.40%    0.67%
85M 21.00% 16.40% 21.80% 27.40% 36.60% 28.00% 26.80% 40.20% 41.80% 37.60% 22.80% 29.13%
300M 46.20% 69.00% 72.80% 78.60% 76.40% 70.40% 71.80% 74.80% 73.80% 78.40% 66.20% 70.76%
Block 85M    5.60%    2.40%    0.80%    0.80%    0.20%    1.00%    0.80%    1.00%    2.60%    1.80%    6.40%    2.13%
300M 23.40% 52.60% 52.60% 46.60% 46.00% 49.20% 58.40% 70.40% 64.00% 53.60% 18.40% 48.65%
800M 35.80% 74.00% 76.40% 78.40% 69.80% 77.40% 76.40% 79.00% 75.20% 72.80% 53.60% 69.89%
1.2B 57.20% 86.60% 88.80% 85.60% 80.40% 85.20% 90.40% 89.20% 91.00% 90.40% 78.80% 83.96%
Table 6: Accuracy on Needle-In-a-Haystack task with Verbatim prompt.
Depth
Models N-Emb 0 10 20 30 40 50 60 70 80 90 100 Mean
Vanilla 19M    8.20%    1.40%    3.00%    6.80%    7.80%    12.60% 45.40% 65.80% 63.40% 84.60% 99.40% 36.22%
85M 95.60% 99.40% 99.00% 99.40% 99.20% 99.20% 99.00% 99.60% 99.60% 99.00% 95.60% 98.60%
300M 99.60% 100.00% 100.00% 99.80% 100.00% 100.00% 99.80% 100.00% 100.00% 100.00% 99.80% 99.91%
Block 85M 96.20% 97.60% 96.20% 96.60% 98.40% 98.00% 97.20% 98.80% 99.00% 99.40% 96.20% 97.60%
300M 90.20% 99.40% 99.60% 99.20% 98.60% 99.60% 99.60% 99.80% 99.80% 99.20% 99.20% 98.56%
800M 95.20% 99.40% 98.80% 98.80% 98.80% 98.80% 99.00% 99.40% 99.20% 97.40% 99.60% 98.58%
1.2B 92.60% 98.40% 99.40% 98.80% 99.60% 99.60% 98.80% 99.80% 99.80% 99.20% 98.00% 98.55%
Appendix P Uptraining strategy for training efficiency
@@
Ainslie et al. [2] have demonstrated the significance of weight initialization for effectively uptraining models. Our extensive ablation studies reveal the optimal strategies for the Block Tramsformers: (1) Dividing a vanilla transformer layer in half and assigning each half to the block and token decoders, respectively, outperforms assigning the same weights of selected layers to both. (2) Initializing the input block embedding as the average of token embeddings within the block improves performance. (3) Initializing token decoder prefixes by replicating the context embedding enhances convergence. As depicted in Figure 21, these initialization techniques allow uptrained models to nearly match fully pretrained models. While larger models generally require longer uptraining, this approach still converges faster and recovers performance better than random initialization.
@@
Figure 21: Training loss curve of uptraining strategy for two model sizes. Scratch denotes pretraining models from randomly initialized weights. The numbers in parentheses represents the number of training tokens.
Appendix Q Performance comparison to MEGABYTE
@@
Appendix R Visualization of attention scores in Block Transformer
@@
Appendix S Analysis on the context block embedding